### v1

# On some practical challenges of conformal prediction

Abstract
Conformal prediction is a model-free machine learning method for creating prediction regions with a guaranteed coverage probability level. However, a data scientist often faces three challenges in practice: (i) the determination of a conformal prediction region is only approximate, jeopardizing the finite-sample validity of prediction, (ii) the computation required could be prohibitively expensive, and (iii) the shape of a conformal prediction region is hard to control. This article offers new insights into the relationship among the monotonicity of the non-conformity measure, the monotonicity of the plausibility function, and the exact determination of a conformal prediction region. Based on these new insights, we propose a simple strategy to alleviate the three challenges simultaneously.
Keywords and phrases: Data science; exact determination of conformal prediction regions; explainable machine learning; finite-sample validity.
1 Introduction
Suppose $Z_{1}=(X_{1},Y_{1}),Z_{2}=(X_{2},Y_{2}),\ldots$ is a sequence of exchangeable random vectors, where $X_{i}\in\mathbb{R}^{p}$ for $p\geq 1$, $Y_{i}\in\mathbb{R}$, and each $Z_{i}$ follows a distribution $\mathsf{P}$. Our goal is to perform an interval prediction of the next response $Y_{n+1}$ at a randomly sampled feature $X_{n+1}$, based on past observations of $Z^{n}=\{Z_{1},\ldots,Z_{n}\}$. If one takes a parametric model approach, there will be two potential dangers lurking behind the scene: (i) model misspecification (e.g., Claeskens and Hjort 2008) and (ii) the effect of selection (e.g., Leeb 2009; Berk et al. 2013; Hong et al. 2018; Kuchibhotal et al. 2022). While these two issues can be largely avoided by employing a nonparametric model, nearly all nonparametric models have tuning parameters and are only asymptotically valid. These issues associated with a model-based approach prompted researchers to seek a model-free approach for creating valid prediction regions. Early works in this direction include Wilks (1941), Fligner and Wolfe (1976), and Frey (2013). However, these works only treat the unsupervised learning case. It is unclear how the methods proposed by these papers can be applied to the regression problem without disregarding the information supplied by predictors. For prediction in the regression setting, conformal prediction (e.g., Vovk et al. 2005, 2009; Shafer and Vovk 2008; Barber et al. 2021) is a model-free machine learning method for generating finite-sample valid prediction regions at a given confidence level.
To apply conformal prediction, we first choose a non-conformity measure $M(B,z)$ which is a real-valued deterministic mapping of two arguments, where the first argument $B=\{z_{1},\ldots,z_{n}\}$ is a bag, i.e., a collection, of observed data and the second argument $z=(x,y)$ is a provisional value of a future observation. Then we run the following Algorithm 1:
1Initialize: data $z^{n}=\{z_{1},\ldots,z_{n}\}$ and $x_{n+1}$, non-conformity measure $M$, and a possible $y$ value;
2 Set $z_{n+1}=(x_{n+1},y)$ and write $z^{n+1}=z^{n}\cup\{z_{n+1}\}$;
3 Define $\mu_{i}=M(z^{n+1}\setminus\{z_{i}\},z_{i})$ for $i=1,\ldots,n,n+1$;
4 Compute $\mathsf{pl}_{x_{n+1},z^{n}}(y)=(n+1)^{-1}\sum_{i=1}^{n+1}1\{\mu_{i}\geq\mu_{n+1}\}$;
5 Return $\mathsf{pl}_{x_{n+1},z^{n}}(y)$;
Algorithm 1 Conformal prediction (supervised learning)
[…]
$C_{\alpha}(x;z^{n})=\{y:\mathsf{pl}_{x_{n+1},Z^{n}}(y)>\alpha\},$ (1)
where $0<\alpha<1$. The basic properties of the rank statistic imply the next theorem.
Theorem 1.
Suppose $Z_{1},Z_{2},\ldots$ is a sequence of exchangeable random vectors and each $Z_{i}$ is generated from a distribution $\mathsf{P}$. Let $\mathsf{P}^{n+1}$ denote the corresponding joint distribution of $Z^{n+1}=\{Z_{1},\ldots,Z_{n},Z_{n+1}\}$. For $\alpha\in(0,1)$, define $t_{n}(\alpha)=(n+1)^{-1}\lfloor(n+1)\alpha\rfloor$, where $\lfloor a\rfloor$ denotes the greatest integer less than or equal to $a$. Then
$\sup\mathsf{P}^{n+1}\{\mathsf{pl}_{X_{n+1},Z^{n}}(Y_{n+1})\leq t_{n}(\alpha)\}\leq\alpha\quad\text{for all $n$ and all $\alpha\in(0,1)$},$
where the supremum is over all distributions $\mathsf{P}$ for $Z_{1}$.
It follows from Theorem 1 that the prediction region given by (1) is finite-sample valid in the sense that
$\mathsf{P}^{n+1}\{Y_{n+1}\in C_{\alpha}(X_{n+1};Z^{n})\}\geq 1-\alpha\quad\text{for all $(n,\mathsf{P})$},$ (2)
where $\mathsf{P}^{n+1}$ is the joint distribution for $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n}),(X_{n+1},Y_{n+1})$. This finite-sample validity says the coverage probability of the conformal prediction region is no less than the advertised confidence level.
While (2) guarantees that the conformal prediction region $C_{\alpha}(x;Z^{n})$ is finite-sample validity for any non-conformity measure $M$, a data scientist faces several challenges in practice. First, it is clear from (1) that exact determination of a conformal prediction region generally requires one to run Algorithm 1 for all possible $y\in\mathbb{R}$. This is practically impossible. This is the reason why most existing papers only implement Algorithm 1 for a grid of $y$ values. Let $\widehat{C}_{\alpha}(Z_{n+1};Z^{n})$ denote the set results from such an approximation. Then the finite-sample validity of $\widehat{C}_{\alpha}(Z_{n+1};Z^{n})$ is nowhere justified. It is important to underline that this challenge is not to be confused with the finite-sample validity of the conformal prediction region $C_{\alpha}(X_{n+1};Z^{n})$. The finite sample validity of $C_{\alpha}(X_{n+1};Z^{n})$ is guaranteed by Theorem 1. But the set $\widehat{C}_{\alpha}(Z_{n+1};Z^{n})$ is not the same as the conformal prediction region $C_{\alpha}(X_{n+1};Z^{n})$. Therefore, one cannot use the finite-sample validity of $C_{\alpha}(X_{n+1};Z^{n})$ to justify the finite-sample validity of $\widehat{C}_{\alpha}(X_{n+1};Z^{n})$. The second challenge is closely related to the first. Even if we choose to consider only a grid of $y$ values, the computation needed for a determining $\widehat{C}_{\alpha}(Z_{n+1};Z^{n})$ could still be prohibitively expensive, although some methods, such as split conformal prediction, have been proposed to circumvent this challenge; see, for example, Lei et al. (2018). Finally, the prediction region $C_{\alpha}(X_{n+1};Z^{n})$ is not guaranteed to be an interval. In general, $C_{\alpha}(Z_{n+1};Z^{n})$ can be a disjoint union of several non-overlapping intervals, which is inappropriate for many applications; see, for example, Lei et al. (2013). For many practical purposes, a data scientist often needs a prediction region to be an interval of a certain shape, such as $(-\infty,a)$, $(a,b)$, or $(b,\infty)$, where $a$ and $b$ are real numbers.
Among these three challenges, the first one is the most serious one because it affects the finite-sample validity of prediction—the key selling point of conformal prediction. However, it is rarely addressed in the literature. Two exceptions are Hong and Martin (2021) and Hong (2025). Hong (2025) finds a suitable non-conformity measure $M$ so that
$\mu_{i}\geq\mu_{n+1}\text{ if and only if $Y_{i}\geq f(X^{n+1})+Y_{n+1}$ or $Y_{i}\leq f(X^{n+1})+Y_{n+1}$,}$ (3)
where $X^{n+1}=\{X_{1},\ldots,X_{n},X_{n+1}\}$ and $f$ is some real-valued function. Then, (3) implies that the corresponding plausibility function is monotonic in $y$, which further implies the resulting conformal prediction region $C_{\alpha}(X_{n+1},Z^{n})$ equals the prediction interval based on some order statistics. The ad hoc strategy employed by Hong and Martin (2021) for unsupervised learning is of the same spirit, though their choice of their non-conformity measure leads to
$\mu_{i}\geq\mu_{n+1}\text{ if and only if $X_{i}\geq X_{n+1}$ or $X_{i}\leq X_{n+1}$},$ (4)
where $X_{i}\in\mathbb{R}$ for $1\leq i\leq n+1$. In both cases, the three aforementioned challenges are overcome simultaneously: the determination of $C_{\alpha}(X_{n+1},Z^{n})$ is exact, the computation needed is simple, and the shape of the $C_{\alpha}(X_{n+1},Z^{n})$ is an interval. In general, if we can determine $C_{\alpha}(X_{n+1},Z^{n})$ exactly, then the computational challenge is likely to be vanish, though care is still needed to ensure $C_{\alpha}(X_{n+1},Z^{n})$ is an interval of a desired form.
Recall that a non-conformity measure $M$ is said to be monotonically increasing if
$y\leq y^{\prime}\Longrightarrow M(B,y)\leq M(B,y^{\prime});$
$M$ is said to be monotonically decreasing if
$y\leq y^{\prime}\Longrightarrow M(B,y)\geq M(B,y^{\prime}).$
We say $M$ is monotonic if it is either monotonically increasing or monotonically decreasing.
Given the above observations regarding Hong and Martin (2021) and Hong (2025), it is natural to ask whether we can say something about the relationship among the monotonicity of $M$, the monotonicity of the plausibility function, Property (3), the shape of the conformal prediction region, and exact determination of the conformal prediction region $C_{\alpha}(X_{n+1},Z^{n})$. In particular, there are several open questions:
  1. (I)
    Does monotonicity of the non-conformity measure $M$ imply (3)?
  2. (II)
    Does (3) imply the monotonicity of $M$?
  3. (III)
    Is monotonicity of $M$ a necessary condition for $C_{\alpha}(X_{n+1},Z^{n})$ to be an interval?
  4. (IV)
    Does the monotonicity of $M$ imply that the resulting conformal prediction region is an interval?
  5. (V)
    Is (3) a necessary condition for $C_{\alpha}(X_{n+1},Z^{n})$ to be an interval? (Note that the converse is true: (3) implies $C_{\alpha}(X_{n+1},Z^{n})$ is a one-sided interval.)
  6. (VI)
    Is monotonicity of the plausibility function $\mathsf{pl}_{x_{n+1},z^{n}}$ a necessary condition for $C_{\alpha}(X_{n+1},Z^{n})$ to be a one-sided interval? (Note that the converse holds, i.e., monotonicity of the plausibility function implies $C_{\alpha}(X_{n+1},Z^{n})$ is a one-sided interval.)
  7. (VII)
    Is (3) a necessary condition for the exact determination of $C_{\alpha}(X_{n+1},Z^{n})$? (Note that the converse is true: (3) implies $C_{\alpha}(X_{n+1},Z^{n})$ is a one-sided interval; hence it implies exact determination of $C_{\alpha}(X_{n+1},Z^{n})$.)
  8. (VIII)
    Is monotonicity of the plausibility function $\mathsf{pl}_{x_{n+1},z^{n}}$ a necessary condition for the exact determination of $C_{\alpha}(X_{n+1},Z^{n})$? (Note that the converse holds, i.e., monotonicity of the plausibility function implies $C_{\alpha}(X_{n+1},Z^{n})$ is a one-sided interval; hence, it implies the exact determination of $C_{\alpha}(X_{n+1},Z^{n})$.)
  9. (IX)
    Is the monotonicity of $M$ a necessary condition for the exact determination of $C_{\alpha}(X_{n+1},Z^{n})$?
In this article, we offer new insights into conformal prediction by answering the above questions. These new insights suggest that it is challenging to find a hard-and-fast rule for choosing a non-conformity measure. In view of this fact and the principle of parsimony, we propose a simple strategy to overcome the aforementioned three challenges simultaneously. The remainder of the paper is organized as follows. Section 2 answers Questions (I)–(IX). Section 3 details our general strategy to overcome the three practical challenges of conformal prediction. Section 4 provides two numerical examples based simulation to demonstrate the excellent performance of the proposed method. Section 5 concludes the article with some remarks. The insights in Section 2 and the strategy in Section 3 have their counterparts in unsupervised learning, which are given in the Appendix.
2 Answers to Questions (I)—(IX)
Here, we answer Questions (I)—(IX). If we answer an open question in the affirmative, we will give a proof; otherwise, we will give a counterexample. Henceforth, we will use the following notation. Let $B=\{z_{1},\ldots,z_{n}\}$ be a bag of observations of size $n$. For $i=1,\ldots,n$, let $z_{i}=(x_{i1},\ldots,x_{ip},y_{i})$ be the $i$-th observation in $B$. That is, $x_{ij}$ denotes the $i$-th observation of the $j$-th feature. $z=(x_{1},\ldots,x_{p},y)$ will denote a provisional value of a future observation to be predicted.
Answer to Question (I): No. Monotonicity of $M$ need not imply (3).
Example 1.
Let $p=1$ and $M(B,z)=(\sum_{j=1}^{n}x_{j1}+x)+\min\{y_{1},\ldots,y_{n}\}+y$. Then $M$ is monotonic. Also,
$\mu_{i}$ $=$ $M(Z^{n+1}\backslash Z_{i},Z_{i})$
$=$ $\sum_{j=1}^{n+1}X_{j1}+\min\{Y_{1},\ldots,Y_{i-1},Y_{i+1},\ldots,Y_{n},Y_{n+1}\}+Y_{i},\quad i=1,\ldots,n,n+1.$
Thus,
$\mu_{i}\geq\mu_{n+1}$ $\Longleftrightarrow$ $\min\{Y_{1},\ldots,Y_{i-1},Y_{i+1},\ldots,Y_{n},Y_{n+1}\}+Y_{i}$ (5)
$\geq\min\{Y_{1},\ldots,Y_{n}\}+Y_{n+1}$
$\Longleftrightarrow$ $\min\{m_{i},Y_{n+1}\}+Y_{i}\geq\min\{m_{i},Y_{i}\}+Y_{n+1},$
where $m_{i}=\min\{Y_{1},\ldots,Y_{i-1},Y_{i+1},\ldots,Y_{n}\}$. If $m_{i}\geq Y_{n+1}$, then the last inequality of (5) becomes
$Y_{n+1}+Y_{i}\geq\min\{m_{i},Y_{i}\}+Y_{n+1},$
which is trivially true. When $m_{i}<Y_{n+1}$, the last inequality of (5) is equivalent to
$m_{i}+Y_{i}\geq\min\{m_{i},Y_{i}\}+Y_{n+1},$
i. e., $m_{i}<Y_{n+1}\leq Y_{i}+m_{i}-\min\{m_{i},Y_{i}\}$. Now if $m_{i}\leq Y_{i}$, then $m_{i}<Y_{n+1}\leq Y_{i}$. However, if $m_{i}>Y_{i}$ we would have $m_{i}<Y_{n+1}\leq m_{i}$, which is absurd. Therefore, (3) does not hold in this case.
Answer to Question (II): No. (3) does not imply $M$ is monotonic.
Example 2.
Let $p=1$ and $M(B,z)=(\sum_{j=1}^{n}x_{j1}+x)+(y_{1}^{2}+\ldots+y_{n}^{2})+y^{2}+y$. Clearly, $M$ is not monotonic. We have
[…]
Hence, $\mu_{i}\geq\mu_{n+1}$ if and only if
$\sum_{j=1,j\neq i}^{n+1}Y_{j}^{2}+Y_{i}^{2}+Y_{i}\geq\sum_{j=1}^{n}Y_{j}^{2}+Y_{n+1}^{2}+Y_{n+1},$
which is equivalent to $Y_{n+1}\leq Y_{i}$. Therefore, (3) holds.
Answer to Question (III): No. Monotonicity of $M$ is not a necessary condition for $C_{\alpha}(X_{n+1},Z^{n})$ to be an interval?
Example 3.
Consider Example 2. $M$ is not monotonic. However, (3) holds, which implies $C_{\alpha}(X_{n+1},Z^{n})=(-\infty,Y_{(k)})$, where $Y_{(k)}$ is the $k$-th order statistic of $Y_{1},\ldots,Y_{n}$.
Answer to Question (IV): No. Monotonicity of $M$ need not imply $C_{\alpha}(X_{n+1},Z^{n})$ is an interval.
Example 4.
Let $p=1$ and $M(B,z)=(\sum_{j=1}^{n}x_{j1}+x)+y_{1}^{2}+\ldots+y_{n}^{2}+y$. Then $M$ is monotonic. We have
$\mu_{i}=M(Z^{n+1}\backslash Z_{i},Z_{i})=\sum_{j=1}^{n+1}X_{j1}+\sum_{j=1,j\neq i}^{n+1}Y_{j}^{2}+Y_{i},\quad i=1,\ldots,n,n+1.$
Thus,
$\mu_{i}\geq\mu_{n+1}\Longleftrightarrow\sum_{j=1,j\neq i}^{n+1}Y_{j}^{2}+Y_{i}\geq\sum_{j=1}^{n}Y_{j}^{2}+Y_{n+1}.$
It follows that
$\mu_{i}\geq\mu_{n+1}\Longleftrightarrow Y_{n+1}^{2}+Y_{i}\geq Y_{i}^{2}+Y_{n+1},$
implying $Y_{n+1}\in(-\infty,\min\{Y_{i},1-Y_{i}\})\cup(\max\{Y_{i},1-Y_{i}\},\infty)$.
Answer to Question (V): No. (3) is not a necessary condition for $C_{\alpha}(X_{n+1},Z^{n})$ to be an interval.
Example 5.
Consider Example 1. We already know that (3) does not hold in this case. Now note that (5) implies that
$C_{\alpha}(X_{n+1},Z^{n})=(-\infty,a_{(k)}),$
where $a_{i}=Y_{i}+\min\{Y_{1},\ldots,Y_{i-1},Y_{i+1},\ldots,Y_{n}\}-\min\{Y_{1},\ldots,Y_{n}\}$. Therefore, $C_{\alpha}(X_{n+1},Z^{n})$ is a one-sided interval.
Answer to Question (VI): Yes. Monotonicity of the plausibility function $\mathsf{pl}_{x_{n+1},z^{n}}$ is a necessary condition for $C_{\alpha}(X_{n+1},Z^{n})$ to be a one-sided interval.
Proof.
We will prove this fact by contradiction. Suppose $C_{\alpha}(X_{n+1},Z^{n})$ is a one-sided interval. Without loss of generality, we assume the plausibility function $\mathsf{pl}_{x_{n+1},z^{n}}$ is not monotonically increasing. Then, there exist three numbers $a<b<c$ such that $\mathsf{pl}_{x_{n+1},z^{n}}(b)>\mathsf{pl}_{x_{n+1},z^{n}}(a)$ and $\mathsf{pl}_{x_{n+1},z^{n}}(b)>\mathsf{pl}_{x_{n+1},z^{n}}(c)$. Thus, for any confidence level $\alpha$ such that $\max\{\mathsf{pl}_{x_{n+1},z^{n}}(a),\mathsf{pl}_{x_{n+1},z^{n}}(c)\}<\alpha<\mathsf{pl}_{x_{n+1},z^{n}}(b)$, we will have $b\in C_{\alpha}(X_{n+1},Z^{n})$ but $a\not\in C_{\alpha}(X^{n})$ and $c\not\in C_{\alpha}(X_{n+1},Z^{n})$. Therefore, $C_{\alpha}(X_{n+1},Z^{n})$ cannot be a one-sided interval. ∎
Answer to Question (VII): No. Monotonicity of $M$ is not a necessary condition for the exact determination of $C_{\alpha}(X_{n+1},Z^{n})$.
Example 6.
Consider Example 2. In this case, $C_{\alpha}(X_{n+1},Z^{n})=(-\infty,Y_{(k)})$. Thus, we can determine $C_{\alpha}(X_{n+1},Z^{n})$ exactly, though $M$ is not monotonic.
Answer to Question (VIII): No. (3) is not a necessary condition for the exact determination of $C_{\alpha}(X_{n+1},Z^{n})$.
Example 7.
Consider Example 5.
Answer to Question (IX): No. Monotonicity of the plausibility function $\mathsf{pl}_{x_{n+1},z^{n}}$ is not a necessary condition for the exact determination of $C_{\alpha}(X_{n+1},Z^{n})$.
Example 8.
Consider Example 4.
Remark. We did not ask the converse of Question (IX), i.e., whether the monotonicity of $M$ implies that $C_{\alpha}(X_{n+1},Z^{n})$ can be determined exactly, because that question seems to be too broad to be well-defined. The next example illustrates this point.
Example 9.
Let $p=1$ and $M(B,z)=\sum_{i=1}^{n+1}x_{i1}+x+e^{(\max\{0,y\})^{8}}+\max\{0,y\}$. Then $M$ is monotonically increasing. In this case,
$\mu_{i}=M(Z^{n+1}\backslash Z_{i},Z_{i})=\sum_{j=1}^{n+1}X_{j1}+e^{(\max\{0,Y_{i}\})^{8}}+\max\{0,Y_{i}\},\quad i=1,\ldots,n,n+1.$
In particular, we have
$\mu_{n+1}=M(Z^{n+1}\backslash Z_{i},Z_{i})=\sum_{j=1}^{n+1}X_{j1}+e^{(\max\{0,Y_{n+1}\})^{8}}+\max\{0,Y_{n+1}\}.$
It follows that $\mu_{i}\geq\mu_{n+1}$ if and only if
$e^{(\max\{0,Y_{n+1}\})^{8}}+\max\{0,Y_{n+1}\}-\left[e^{(\max\{0,Y_{i}\})^{8}}+\max\{0,Y_{i}\}\right]\leq 0,$
which is not known to have any closed-formula solutions. Therefore, we do not know any method for determining $C_{\alpha}(X_{n+1},Z^{n})$ exactly. This does not mean we cannot find such a method in the future, nor does it imply that such a method does not exist.
3 Proposed strategy
First, we make a simple but important observation.
Theorem 2.
If $\lfloor(n+1)\alpha\rfloor\leq 1$, then the $(1-\alpha)\%$ prediction region $C_{\alpha}(x;Z^{n})$ given by (1) is $\mathbb{R}$.
Proof.
We discuss two possible cases: (i) $\lfloor(n+1)\alpha\rfloor<1$ and (ii) $\lfloor(n+1)\alpha\rfloor=1$. We have $Y_{n+1}\in C_{\alpha}(X_{n+},Z^{n})$ if and only if $\mathsf{pl}_{X_{n+1},Z^{n}}(Y_{n+1})>\lfloor(n+1)\alpha\rfloor/(n+1)$ if and only if $\sum_{i=1}^{n+1}1_{\{\mu_{i}\geq\mu_{n+1}\}}>\lfloor(n+1)\alpha\rfloor$. Since $1_{\{\mu_{n+1}\geq\mu_{n+1}\}}=1$, we have $C_{\alpha}(X_{n+1},Z^{n})=\mathbb{R}$ in Case (i). In Case (ii), we have $1/(n+1)\leq\alpha<2/(n+1)$ and $t_{n}(\alpha)=1/(n+1)$. Since $\mu_{1},\ldots,\mu_{n+1}$ are exchangeable, $\mathsf{pl}_{X_{n+1},Z^{n}}(Y_{n+1})$ follows the discrete uniform distribution on the set $\{1/(n+1),2/(n+1),\ldots,n/(n+1),1\}$. Therefore,
$\mathsf{P}^{n+1}\{\mathsf{pl}_{X_{n+1},Z^{n}}(Y_{n+1})>t_{n}(\alpha)\}=\frac{n}{n+1}\geq 1-\alpha,\quad\text{for all $Y_{n+1}$}.$
It follows that $C_{\alpha}(X_{n+1},Z^{n})=\mathbb{R}$. ∎
Thus, if we want a nontrivial conformal prediction region (i.e., $C_{\alpha}(x;Z^{n})\neq\mathbb{R}$), we must require $\lfloor(n+1)\alpha\rfloor\geq 2$. For the remainder of this article, we let $r_{1}=\min\{n,\lfloor(n+1)(1-\alpha)\rfloor+1\}$, $r_{2}=(n+1)-\lfloor(n+1)(1-\alpha)\rfloor$, and $r_{3}=(n+1)-\lfloor(n+1)\alpha\rfloor$.
The negative answers to most questions in the previous section show that it is challenging to give a hard-and-fast rule for choosing a non-conformity measure so that the resulting conformal prediction region can be determined exactly and can be of the desired shape. In the extant literature, many existing statistical models, such as ordinary linear regression, ridge regression, lasso, and kernel density estimation, have been used to create non-conformity measures; see Vovk et al. (2005) and Lei et al. (2014), and references therein. However, barely any work discusses how to determine exactly the corresponding conformal prediction regions based on these non-conformity measures; the numerical examples based on these proposed non-conformity measures are mostly based on approximate determination of the corresponding conformal prediction regions. As pointed out in Section 2, the resulting regions have no provable finite-sample validity. Since conformal prediction is a model-free method, one does not have to use a complicated statistical model to create a non-conformity measure. There is no evidence, either theoretical or practical, that doing so would have any advantage. It is evident that if (3) holds, then we can address the three practical challenges simultaneously. There are numerous choices of the non-conformity measures that can lead to (3). Driven by the principle of parsimony, we prefer something simple, such as a linear polynomial in data. However, to obtain a two-sided/bounded prediction interval, we will need something other than (3). In fact, a multivariate polynomial of data with degree two suffices, as we will see below. Therefore, we propose the following non-conformity measure.
For $i=1,\ldots,n$, we write $X_{i}$ as $(X_{i1},\ldots,X_{ip})$ where each $X_{ij}\in\mathbb{R}$. That is, $X_{ij}$ denotes the $i$-th observation of the $j$-th predictor. Hence, $(X_{i1},\ldots,X_{ip},Y_{i})=(X_{i},Y_{i})$ is the $i$-th observation. Suppose $B=\{(x_{11},\ldots,x_{1p},y_{1}),\ldots,(x_{n1},\ldots,x_{np},y_{n})\}$ and $z=(x_{1},\ldots,x_{p},y)$ is a provisional value of $Z_{n+1}$. Consider the following nonconformity measure
$M(B,z)=(\beta_{2}y^{2}+\beta_{1}y)+\gamma\left[\sum_{j=1}^{p}x_{j}+\eta\sum_{i=1}^{n}\left(y_{i}-\sum_{j=1}^{p}x_{ij}\right)\right],$ (6)
where $\beta_{1},\beta_{2},\gamma$, and $\eta$ are constants to be chosen at the discretion of the data scientist.
The next three theorems show that this non-conformity measure has several advantages. First, the aforementioned three practical challenges will be addressed simultaneously. In particular, we can determine the resulting conformal prediction regions exactly. Second, it can generate conformal prediction intervals of three different shapes. Indeed, $\beta_{2},\beta_{1}$, and $\gamma$ control the shape of the conformal prediction regions. For example, if $\beta_{2}=0$ and $\beta_{1}=1$, the resulting conformal prediction region will be a one-sided interval; when $\beta_{2}\neq 0$ and $\beta_{1}=0$, the resulting conformal prediction region will be a bounded (two-sided) interval.
We do not use a higher-order polynomial because such a choice will involve more complicated computation, and will likely lead to a challenging case in determining the conformal prediction regions, similar to what we saw in Example 9. In the next three theorems, we take the coefficients of the second-order polynomial in (6) to be some simple numbers, such as $0,1$, and 1-1. Tedious algebra shows that using other coefficients will not provide any advantages.
Theorem 3.
Suppose $0<\alpha<1$, $\lfloor(n+1)\alpha\rfloor\geq 2$, and the non-conformity measure is given by (6). If $\beta_{2}=0$, $\beta_{1}=1$, $\gamma=-1$, and $1+\eta>0$, then the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n+1};Z^{n})$ is the one-sided interval $(-\infty,a_{(r_{1})})$, where $a_{i}=\sum_{j=1}^{p}(X_{(n+1)j}-X_{ij})+Y_{i}$ for $1\leq i\leq n$ and $a_{(k)}$ is the $k$-th ordered value of $a_{1},\ldots,a_{n}$.
Proof.
For $i=1,\ldots,n,n+1$, let $S_{i}$ denote the sum $\sum_{j=1}^{p}X_{ij}$. Then
$\mu_{i}$ $=$ $M(Z^{n+1}\backslash\{Z_{i}\},Z_{i})=Y_{i}-\left[S_{i}+\eta\sum_{j=1,j\neq i}^{n+1}(Y_{j}-S_{j})\right],\ i=1,\ldots,n,$
$\mu_{n+1}$ $=$ $M(Z^{n},Z_{i})=Y_{n+1}-\left[S_{n+1}+\eta\sum_{j=1}^{n}(Y_{j}-S_{j})\right].$
Therefore, $\mu_{i}\geq\mu_{n+1}$ if and only if
$S_{i}+\eta\sum_{j=1,j\neq i}^{n+1}(Y_{j}-S_{j})-Y_{i}\leq S_{n+1}+\eta\sum_{j=1}^{n}(Y_{j}-S_{j})-Y_{n+1},$
which is equivalent to
$S_{i}-Y_{i}+\eta\sum_{j=1}^{n}(Y_{j}-S_{j})-\eta(Y_{i}-S_{i})+\eta(Y_{n+1}-S_{n+1})\leq S_{n+1}-Y_{n+1}+\eta\sum_{j=1}^{n}(Y_{j}-S_{j}).$
Since $1+\eta>0$, the last display implies $\mu_{i}\geq\mu_{n+1}$ if and only if $Y_{n+1}\leq(S_{n+1}-S_{i})+Y_{i}$ for $1\leq i\leq n$. Therefore, the theorem follows from (1) and the definition of the plausibility function $\mathsf{pl}_{x_{n+1},z^{n}}$. ∎
Theorem 4.
Suppose $0<\alpha<1$, $\lfloor(n+1)\alpha\rfloor\geq 2$, and the non-conformity measure is given by (6). If $\beta_{2}=0$, $\beta_{1}=-1$, $\gamma=1$, and $1+\eta<0$ and $\eta\neq-1$,, then the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n+1};Z^{n})$ is the one-sided interval $(a_{(r_{2})},\infty)$, where $a_{i}=\sum_{j=1}^{p}(X_{(n+1)j}-X_{ij})+Y_{i}$ for $1\leq i\leq n$.
Proof.
The proof is completely similar to that of Theorem 3. ∎
Theorem 5.
Suppose $0<\alpha<1$, $\lfloor(n+1)\alpha\rfloor\geq 2$, and the non-conformity measure is given by (6). If $\beta_{2}=1$, $\beta_{1}=0$, $\gamma=-1$, and $\eta\geq\max_{1\leq i\leq n}\{2(\sqrt{\max\{0,c_{i}^{2}-d_{i}\}}-c_{i})\}$ where $c_{i}=Y_{i}+\sum_{j=1}^{p}(X_{(n+1)j}-X_{ij})$ and $d_{i}=Y_{i}^{2}+\sum_{j=1}^{p}(X_{(n+1)j}-X_{ij})$ for $1\leq i\leq n$, then the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n+1};Z^{n})$ is the bounded interval $(a_{(r_{1})},b_{(r_{3})})$, where $a_{i}=-\sqrt{\eta^{2}/4+\eta c_{i}+d_{i}}-\eta/2$ and $b_{i}=\sqrt{\eta^{2}/4+\eta c_{i}+d_{i}}-\eta/2$ for $1\leq i\leq n$.
Proof.
We still let $S_{i}$ denote the sum $\sum_{j=1}^{p}X_{ij}$ for $i=1,\ldots,n,n+1$. Then
$\mu_{i}$ $=$ $Y_{i}^{2}-\left[S_{i}+\eta\sum_{j=1,j\neq i}^{n+1}(Y_{i}-S_{i})\right],\ i=1,\ldots,n,$
$\mu_{n+1}$ $=$ $Y_{n+1}^{2}-\left[S_{n+1}+\eta\sum_{j=1}^{n}(Y_{i}-S_{i})\right].$
Thus, $\mu_{i}\geq\mu_{n+1}$ if and only if
$Y_{i}^{2}-\left[S_{i}+\eta\sum_{j=1}^{n}(Y_{i}-S_{i})-\eta(Y_{i}-S_{i})+\eta(Y_{n+1}-S_{n+1})\right]\geq Y_{n+1}^{2}-\left[S_{n+1}+\eta\sum_{j=1}^{n}(Y_{i}-S_{i})\right],$
which implies $\mu_{i}\geq\mu_{n+1}$ if and only if
$(Y_{n+1}+\eta/2)^{2}\leq\eta^{2}/4+\eta c_{i}+d_{i}.$ (7)
By the assumption $\eta\geq 2(\sqrt{\max\{0,c_{i}^{2}-d_{i}\}}-c_{i})$, we know the right-hand side of (7) is positive. Therefore, $\mu_{i}\geq\mu_{n+1}$ if and only if
$Y_{n+1}\in I_{i}=(-\sqrt{\eta^{2}/4+\eta c_{i}+d_{i}}-\eta/2,\sqrt{\eta^{2}/4+\eta c_{i}+d_{i}}-\eta/2)=(a_{i},b_{i}),\ i=1,\ldots,n.$
The same argument in the proof of Theorem 2 shows that $Y_{n+1}\in C_{\alpha}(X_{n+1},Z^{n})$ if and only if $\sum_{i=1}^{n}1_{\{\mu_{i}\geq\mu_{n+1}\}}>r=\lfloor(n+1)\alpha\rfloor$. Note that $Y_{n+1}\not\in(a_{(r_{1})},b_{(r_{3})})$ if and only if $Y_{n+1}$ does not belong to at most r1r-1 $I_{i}$’s, or equivalently, $Y_{n+1}\in(a_{(r_{1})},b_{(r_{3})})$ if and only if $Y_{n+1}$ belongs to at least $r$ $I_{i}$’s. Therefore, $Y_{n+1}\in C_{\alpha}(X_{n+1},Z^{n})$ if and only if $Y_{n+1}\in(a_{(r_{1})},b_{(r_{3})})$. ∎
4 Illustration
Throughout this section, we will let ${\sf N}(\mu,\sigma^{2})$ denote the normal distribution with mean $\mu$ and variance $\sigma^{2}$, and let ${\sf Unif}(a,b)$ denote the uniform distribution supported on $(a,b)$.
Recall the standard linear regression model
$Y=X\beta+\sigma\varepsilon,$ (8)
where $Y$ is a $n$-dimensional vector of response variables, $X$, called data/design matrix, is an $n\times p$ matrix of observations of predictors, $\varepsilon$ is an $n$-dimensional vector of iid standard normal errors, $\beta$ is a $p$-dimensional vector of regression coefficients, and $\sigma>0$ is the scale parameter; if the model includes an intercept term, then the first column of $X$ consists of a $n$-vector of 1s and $X$ will be an $n\times(p+1)$ matrix. If we fit this model with the least squares estimation, the $100(1-\alpha)$% prediction interval for $Y_{n+1}$ at $X_{n+1}=x$, is
$x^{T}\hat{\beta}\pm t_{n-p}(\alpha/2)\,\hat{\sigma}\,\{1+x^{\top}(X^{\top}X)^{-1}x\}^{1/2},$ (9)
where $\hat{\beta}$ is the least squares estimator of $\beta$, $\hat{\sigma}$ is the residual standard error, $t_{\nu}(\alpha)$ denotes the $(1-\alpha)^{\text{th}}$ quantile of the (central) Student-t distribution with $\nu$ degrees of freedom. For $0<\alpha<1$, we will use $q_{\alpha}$ to denote the $(1-\alpha)^{\text{th}}$ quantile of $Y$.
Example A
For $\alpha=0.1$, we generate $N=5,000$ random samples of size $n=1,001$ from the following model:
$Y=X_{1}+X_{2}+\epsilon,$
where $X_{1},$ $X_{2}$, and $\epsilon$ are independent, and $X_{1}\sim{\sf N}(0,2)$, $X_{2}\sim{\sf N}(0,1)$, and $\epsilon\sim{\sf N}(0,\sqrt{0.2})$. For each sample, the response values of the first $1,000$ sample points and all the $1,001$ values of the two features are used to construct the conformal prediction intervals in the above three theorems as well as the linear model prediction intervals of the same three shapes. The two prediction intervals in Theorem 3 and Theorem 4 are independent of $\eta$. For the bounded prediction interval in Theorem 5, we take $\eta=\max_{1\leq i\leq n}\{2(\sqrt{\max\{0,c_{i}^{2}-d_{i}\}+1}-c_{i})\}$. Then, the 1,0011,001st response value is treated as the future response value we want to predict. We estimate the coverage probability of each prediction interval as $K/N$ where $K$ is the number of times it contains the 1,0011,001st response value. For the bounded (i.e., two-sided) $100(1-\alpha)\%$ prediction interval, we also calculate the ratio of its expected length to the length of the oracle interval $(q_{1-\alpha/2}-q_{\alpha/2})$. Table 1 summarizes the results.
Prediction interval form Linear Model Conformal Prediction
$\quad\quad(-\infty,a)$ 0.9038 0.9062
$\quad\quad(a,\infty)$ 0.9052 0.9034
$\quad\quad(a,b)$     0.9050 (0.1187)     0.9080 (1.1922)
Table 1: Coverage probabilities (ratio of interval length, if applicable) of the $90\%$ prediction intervals in Example A, based on the linear model and conformal prediction (Theorem 3, Theorem 4, and Theorem 5).
All three types of linear model prediction intervals and conformal prediction intervals achieve the nominal coverage probability. For conformal prediction intervals, this is no surprise because conformal prediction intervals are provably finite-sample valid. Since the linear model is well-specified in this example, the excellent performance of the linear model prediction intervals is expected. The bounded linear model prediction interval given by (9) is more efficient than the bound conformal prediction interval, though both are efficient. Since the conformal prediction interval is distribution-free, it is expected to be more conservative than the linear model prediction interval when the linear model is correct. Note that the oracle interval is based on information of $Y$ only, but the linear model prediction interval, given by (9), is constructed using information from both the response variable and predictors. Intuitively, when the variance of the noise (i.e, the error term $\epsilon$) is dominated by the variance of the predictors (Here $\mathsf{V}(X_{1}+X_{2})=3$ is much larger than $0.2=\mathsf{V}(\epsilon)$.), the oracle interval will be less efficient than the linear model prediction interval.
Example B
We perform the same simulation with the same values of $\alpha$, $N$ and $n$ as in Example A except that data are generated from the following model:
$Y=X_{1}+X_{2}+\epsilon,$
where $X_{1},X_{2}$, and $\epsilon$ are independent and $X_{1}\sim{\sf N}(0,2)$, $X_{2}\sim{\sf N}(0,1)$, and $\epsilon\sim{\sf Unif}(-0.6,0.6)$. The results are summarized in Table 2.
Prediction interval form Linear Model Conformal Prediction
$\quad\quad(-\infty,a)$ 0.8750 0.9052
$\quad\quad(a,\infty)$ 0.8742 0.9034
$\quad\quad(a,b)$     0.9559 (0.1987)     0.9150 (1.1015)
Table 2: Coverage probabilities (ratior of interval length, if applicable) of the $90\%$ prediction intervals in Example B, based on the linear model and conformal prediction (Theorem 3, Theorem 4, and Theorem 5).
In this case, the linear model is incorrect. Therefore, the two one-sided linear model prediction intervals fail to provide adequate coverage. The bounded linear model prediction interval is generally not expected to provide adequate coverage when the linear model is wrong. It achieves the nominal coverage probability here simply because the probability mass of the distribution of the error term $\epsilon$ is concentrated in the interval (0.6,0.6)(-0.6,0.6). The three conformal prediction intervals all achieve the nominal coverage probability, as anticipated. While the bounded conformal prediction interval is a bit conservative, it is efficient when compared to the oracle prediction interval.
5 Concluding remarks
Conformal prediction is a powerful general strategy for creating finite-sample validity prediction intervals. For any non-conformity measure, the corresponding conformal prediction region is guaranteed to be finite-sample valid. However, the determination of a conformal prediction region generally requires a data scientist to evaluate the plausibility function for infinitely many values, which cannot be accomplished in practice. In the prior literature, many authors approximated the conformal prediction regions by evaluating the plausibility function for only a grid of possible values. The resulting region, not the same as the conformal prediction region, no longer has provable finite-sample validity, let alone two other challenges: (i) the computation required can still be prohibitively expensive, and (ii) the resulting prediction region might not be of a desired shape. While confronted with these practical challenges, some colleagues seem to believe that a monotonic nonconformity measure will resolve these issues. Our investigation showed that this is false, among other insights into the relationship between the monotonicity of the non-conformity measures, the monotonicity of the plausibility function, the shape of the conformal prediction region, and the exact determination of the conformal prediction regions. Our investigation also showed that it is challenging to give a hard-and-fast rule for choosing a non-conformity measure so that this issue can be avoided and the resulting conformal prediction region can be of a desired shape.
Driven by the principle of parsimony, we propose a non-conformity measure based on a multivariate polynomial of degree two. When we use the proposed non-conformity measure, we can not only avoid three common practical challenges in conformal prediction but also generate conformal prediction intervals of three common shapes. Of course, our proposal is by no means the only method a data scientist can employ to address the three common practical challenges. But it is simple and easy to implement. However, it does not work in higher-dimension cases. For example, if $Y\in\mathbb{R}^{p}$ for $p\geq 2$, our proposal may not apply, since there is no natural order on $\mathbb{R}^{p}$ for $p\geq 2$.
6 Appendix
This appendix documents the key results in the unsupervised learning.
6.1 Notation and setup
A similar strategy exists for the unsupervised learning setting. Suppose our observations consist of a sequence of exchangeable random variables $X_{1},X_{2},\ldots$, where each $X_{i}$ follows a distribution $\mathsf{P}$. We want to create a prediction interval of the next observation $X_{n+1}$, based on past observations of $X^{n}=\{X_{1},\ldots,X_{n}\}$. The conformal prediction algorithm in this case is the following Algorithm 2:
1Initialize: data $x^{n}=\{x_{1},\ldots,x_{n}\}$ , non-conformity measure $M$;
2 for each possible $x$ value do
3   Set $x_{n+1}=x$ and write $x^{n+1}=x^{n}\cup\{x_{n+1}\}$;
4   Define $\mu_{i}=M(x^{n+1}\setminus\{x_{i}\},x_{i})$ for $i=1,\ldots,n,n+1$;
5   Compute $\mathsf{pl}_{x^{n}}(x)=(n+1)^{-1}\sum_{i=1}^{n+1}1\{\mu_{i}\geq\mu_{n+1}\}$;
6
7 end for
8Return $\mathsf{pl}_{z^{n}}(x)$ for each $x$;
Algorithm 2 Conformal prediction (unsupervised)
A $100(1-\alpha)\%$ conformal prediction region can be constructed as follows:
$C_{\alpha}(X^{n})=\{x:\mathsf{pl}_{Z^{n}}(x)>\alpha\},$ (10)
where $0<\alpha<1$. The finite-sample validity still holds:
$\mathsf{P}^{n+1}\{X_{n+1}\in C_{\alpha}(X_{n+1})\}\geq 1-\alpha\quad\text{for all $(n,\mathsf{P})$},$
where $\mathsf{P}^{n+1}$ is the joint distribution for $(X_{1},\ldots,X_{n},X_{n+1})$.
Let $B=\{x_{1},\ldots,x_{n}\}$ be a bag of observations of size $n$. For $i=1,\ldots,n$, let $(x_{1}\ldots,x_{n})$ be the observations in $B$. Suppose $x$ is a provisional value of a future observation of $X$ to be predicted. In this case, a non-conformity measure $M$ is said to be monotonically increasing if
$x\leq x^{\prime}\Longrightarrow M(B,x)\leq M(B,x^{\prime});$
$M$ is said to be monotonically decreasing if
$x\leq x^{\prime}\Longrightarrow M(B,x)\geq M(B,x^{\prime}).$
$M$ is said to be monotonic if it is either monotonically increasing or monotonically decreasing.
In the unsupervised learning case, (3) corresponds to
$\mu_{i}\geq\mu_{n+1}\text{ if and only if $X_{i}\geq X_{n+1}$ or $X_{i}\leq X_{n+1}$}.$ (11)
Corresponding to Questions (I)—(IX), we have the following Questions (I’)—(IX’).
  1. (I’)
    Does monotonicity of the non-conformity measure $M$ imply (11)?
  2. (II’)
    Does (11) imply the monotonicity of $M$?
  3. (III’)
    Is monotonicity of $M$ a necessary condition for $C_{\alpha}(X^{n})$ to be an interval?
  4. (IV’)
    Does the monotonicity of $M$ imply that the resulting conformal prediction region is an interval?
  5. (V’)
    Is (11) a necessary condition for $C_{\alpha}(X^{n})$ to be an interval? (Note that the converse is true: (11) implies $C_{\alpha}(X^{n})$ is a one-sided interval).
  6. (VI’)
    Is monotonicity of the plausibility function $\mathsf{pl}_{x^{n}}$ a necessary condition for $C_{\alpha}(X^{n})$ to be a one-sided interval? (Note that the converse holds, i.e., monotonicity of the plausibility function implies $C_{\alpha}(X^{n})$ is a one-sided interval.)
  7. (VII’)
    Is the monotonicity of $M$ a necessary condition for the exact determination of $C_{\alpha}(Z^{n})$?
  8. (VIII’)
    Is (11) a necessary condition for exact determination of $C_{\alpha}(Z^{n})$? (Note that the converse is true: (11) implies $C_{\alpha}(Z^{n})$ is a one-sided interval; hence it implies the exact determination of $C_{\alpha}(Z^{n})$.)
  9. (IX’)
    Is monotonicity of the plausibility function $\mathsf{pl}_{z^{n}}$ a necessary condition for the exact determination of $C_{\alpha}(Z^{n})$? (Note that the converse holds, i.e., monotonicity of the plausibility function implies $C_{\alpha}(X^{n})$ is a one-sided interval; hence, it implies the exact determination of $C_{\alpha}(X^{n})$.)
6.2 Answers to Questions (I’)—(IX’)
Answers to Questions (I’)—(IX’) follow from answers to Questions (I)—(IX) since we can simply ignore the predictors and treat the response variable as the random variable of interest in unsupervised learning. For the sake of completeness, we give the details below.
Answer to Question (I’): No. Monotonicity of $M$ need not imply (11).
Example 10.
Let $M(B,x)=\min\{x_{1},\ldots,x_{n}\}+x$. Then $M$ is monotonic. Also,
$\mu_{i}=M(X^{n+1}\backslash X_{i},X_{i})=\min\{X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{n},X_{n+1}\}+X_{i},\quad i=1,\ldots,n,n+1.$
In particular, $\mu_{n+1}=\min\{X_{1},\ldots,X_{n}\}+X_{n+1}$. Thus,
$\mu_{i}\geq\mu_{n+1}$ $\Longleftrightarrow$ $\min\{X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{n},X_{n+1}\}+X_{i}$ (12)
$\geq\min\{X_{1},\ldots,X_{n}\}+X_{n+1}$
$\Longleftrightarrow$ $\min\{m_{i},X_{n+1}\}+X_{i}\geq\min\{m_{i},X_{i}\}+X_{n+1},$
where $m_{i}=\min\{X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{n}\}$. If $m_{i}\geq X_{n+1}$, the last inequiality of (12) becomes
$X_{n+1}+X_{i}\geq\min\{m_{i},X_{i}\}+X_{n+1},$
which is obviously true. When $m_{i}<X_{n+1}$, (12) is equivalent to
$m_{i}+X_{i}\geq\min\{m_{i},X_{i}\}+X_{n+1},$
i. e., $m_{i}<X_{n+1}\leq X_{i}+m_{i}-\min\{m_{i},X_{i}\}$. If $m_{i}\leq X_{i}$, this becomes $m_{i}<X_{n+1}\leq X_{i}$. We cannot have $m_{i}>X_{i}$ because it would imply $m_{i}<X_{n+1}\leq m_{i}$, which is absurd. Therefore, (11) does not hold in this case.
Answer to Question (II’): No. (11) does not imply $M$ is monotonic.
Example 11.
Let $M(B,x)=(x_{1}^{2}+\ldots+x_{n}^{2})+x^{2}+x$. Clearly, $M$ is not monotonic. We have
$\mu_{i}=M(X^{n+1}\backslash\{X_{i}\},X_{i})=\sum_{j=1,j\neq i}^{n+1}X_{j}^{2}+X_{i}^{2}+X_{i},\quad j=1,\ldots,n,n+1.$
Hence, $\mu_{i}\geq\mu_{n+1}$ if and only if
$\sum_{j=1,j\neq i}^{n+1}X_{j}^{2}+X_{i}^{2}+X_{i}\geq\sum_{j=1}^{n}X_{j}^{2}+X_{n+1}^{2}+X_{n+1},$
which is equivalent to $X_{n+1}\leq X_{i}$.
Answer to Question (III’): No. Monotonicity of $M$ is not a necessary condition for $C_{\alpha}(X^{n})$ to be an interval?
Example 12.
Consider Example 11. Here $M$ is not monotonic. However, (11) holds, which implies $C_{\alpha}(X^{n})=(-\infty,X_{(k)})$.
Answer to Question (IV’): No. Monotonicity of $M$ need not imply $C_{\alpha}(X^{n})$ is an interval.
Example 13.
Let $M(B,x)=x_{1}^{2}+\ldots+x_{n}^{2}+x$. Then $M$ is monotonic. We have
$\mu_{i}=M(X^{n+1}\backslash\{X_{i}\},X_{i})=\sum_{j=1,j\neq i}^{n+1}X_{j}^{2}+X_{i}\quad i=1,\ldots,n,n+1.$
Thus,
$\mu_{i}\geq\mu_{n+1}\Longleftrightarrow\sum_{j=1,j\neq i}^{n+1}X_{j}^{2}+X_{i}\geq\sum_{j=1}^{n}X_{j}^{2}+X_{n+1}.$
It follows that
$\mu_{i}\geq\mu_{n+1}\Longleftrightarrow X_{n+1}^{2}+X_{i}\geq X_{i}^{2}+X_{n+1},$
implying $X_{n+1}\in(-\infty,\min\{X_{i},1-X_{i}\})\cup(\max\{X_{i},1-X_{i}\},\infty)$.
Answer to Question (V’): No. (11) is not a necessary condition for $C_{\alpha}(X^{n})$ to be an interval.
Example 14.
Consider Example 10. We know that (11) does not hold in this case. Now note that (12) implies that
$C_{\alpha}(X^{n})=(-\infty,a_{(k)}),$
where $a_{i}=X_{i}+\min\{X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{n}\}-\min\{X_{1},\ldots,X_{n}\}$. Therefore, $C_{\alpha}(X^{n})$ is a one-sided interval.
Answer to Question (VI’): Yes. Monotonicity of the plausibility function $\mathsf{pl}_{x^{n}}$ a necessary condition for $C_{\alpha}(X^{n})$ to be a one-sided interval.
Proof.
Suppose $C_{\alpha}(X^{n})$ is a one-sided interval. Without loss of generality, we assume the plausibility function $\mathsf{pl}_{x^{n}}$ is not monotonically increasing. Then, there are three numbers $a<b<c$ such that $\mathsf{pl}_{x^{n}}(b)>\mathsf{pl}_{x^{n}}(a)$ and $\mathsf{pl}_{x^{n}}(b)>\mathsf{pl}_{x^{n}}(c)$. Thus, for any confidence level $\alpha$ such that $\max\{\mathsf{pl}_{x^{n}}(a),\mathsf{pl}_{x^{n}}(c)\}<\alpha<\mathsf{pl}_{x^{n}}(b)$, we will have $b\in C_{\alpha}(X^{n})$ but $a\not\in C_{\alpha}(X^{n})$ and $c\not\in C_{\alpha}(X^{n})$. Therefore, $C_{\alpha}(X^{n})$ cannot be a one-sided interval. ∎
Answer to Question (VII’): No. Monotonicity of $M$ is not necessary condition for exact determination of $C_{\alpha}(X^{n})$.
Example 15.
Consider Example 11. Here $C_{\alpha}(X^{n})=(-\infty,X_{(k)})$. Hence, we can determine $C_{\alpha}(Z^{n})$ exactly, though $M$ is not monotonic.
Answer to Question (VIII’): No. (11) is not a necessary condition for exact determination of $C_{\alpha}(X^{n})$.
Example 16.
Consider Example 14.
Answer to Question (IX’): No. Monotonicity of the plausibility function $\mathsf{pl}_{z^{n}}$ is not a necessary condition for exact determination of $C_{\alpha}(X^{n})$.
Example 17.
Consider Example 13.
6.3 Proposed strategy
Theorem 6.
If $\lfloor(n+1)\alpha\rfloor\leq 1$, then the $(1-\alpha)\%$ prediction region $C_{\alpha}(X^{n})$ given by (10) is $\mathbb{R}$.
Proof.
Completely similar to the proof of Theorem 2. ∎
Now consider the non-conformity measure
$M(B,x)=\lambda x^{2}+\theta x+\kappa\sum_{j=1}^{n}x_{j},$ (13)
where $B=\{x_{1},\ldots,x_{n}\}$ and $\lambda,\theta$, and $\kappa$ are constants to be decided by the user. Following the same line of reasoning as in the previous section, we can see that the following three theorems hold. Note that Theorems 7 and 8 recover the traditional non-parametric one-sided prediction intervals based on order statistics (e.g., Wiks 941; Fligner and Wolfe 1976; Frey 2013).
Theorem 7.
Suppose $0<\alpha<1$, $\lfloor(n+1)\alpha\rfloor\geq 2$, and the non-conformity measure is given by (13). If $\lambda=0$, $\theta=1$, $\kappa=-1$, then the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n})$ is the one-sided interval $(-\infty,X_{(r_{1})})$.
Theorem 8.
Suppose $0<\alpha<1$, $\lfloor(n+1)\alpha\rfloor\geq 2$, and the non-conformity measure is given by (13). If $\lambda=0$, $\theta=-1$, $\kappa=1$, then the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n})$ is the one-sided interval $(X_{(r_{2})},\infty)$.
Theorem 9.
Suppose $0<\alpha<1$, $\lfloor(n+1)\alpha\rfloor\geq 2$, and the non-conformity measure is given by (13). If $\lambda=1$, $\theta=0$, $\kappa\neq 0$, then the $100(1-\alpha)\%$ conformal prediction region $C_{\alpha}(X_{n})$ is the bounded interval $(a_{(r_{1})},b_{(r_{2})})$, where $a_{i}=\min\{-(\kappa+X_{i}),X_{i}\}$ and $b_{i}=\max\{-(\kappa+X_{i}),X_{i}\}$ for $1\leq i\leq n$.